the prouduct of -15 (and) y is less than or equal to 90 and hi i'm asia i'm fine and how are you to day

Answer by checkley77(12569)

the prouduct of -15 (and) y is less than or equal to 90
-15y<=90
y>=90/-15 [Note: when dividing by a negative value the < sign changes to > sign.
y>=-6 ans.
Proof:
Let y=-5.9
-15*-5.9<=90
88.8<=90